What mass of Br2 is needed to react with 27 g of aluminum? Round to two decimal places.



2Al + 3Br2 → 2AlBr3
1
Expert's answer
2019-03-08T05:50:39-0500

The balanced equation for reaction of aluminium with bromine can be written as follows

"2Al+3Br_{2} \rightarrow 2AlBr_{3}"

"\frac{m(Al)}{N(Al) \cdot M(Al)}=\frac{m(Br_{2})}{N(Br_{2}) \cdot M(Br_{2})}"

"m(Br_{2})=\frac{m(Al) \cdot N(Br_{2}) \cdot M(Br_{2})}{N(Al) \cdot M(Al)}"

"M(Al)=1 \cdot 26.98=26.98 \; amu"

"M(Br_{2})=2 \cdot 79.9=159.8 \; amu"

"m(Br_{2})=\frac{27 \cdot 3 \cdot 159.8}{2 \cdot 26.98} \approx 239.88 \; g"


Answer:

239.88 g